OpenStudy (anonymous):

The sum of five times a number and twice a second number is 18, but the difference of three times the first number and seven times the second number is -22. Find the two numbers using the elimination method. A. -2 and -4 B. -2 and 4 C. 2 and -4 D. 2 and 4

OpenStudy (texaschic101):

5x + 2y = 18 --->(3) 3x - 7y = -22 --->(-5) ------------------ 15x + 6y = 54 (result of multiplying by 3) -15x + 35y = 110 (result of multiplying by -5) -----------------add 41y = 164 y = 4 now sub 4 in for y in either of the original equations 5x + 2y = 18 5x + 2(4) = 18 5x + 8 = 18 5x = 18 - 8 5x = 10 x = 2 check... 3x - 7y = -22 3(2) - 7(4) = -22 6 - 28 = -22 -22 = -22 (correct) x = 2 and y = 4